OpenStudy (anonymous):

Plz help.chose the point - slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2. y-6 =1/2x+3) y=1/2x-6 y+3=1/2 (x-6) x-2y =12

OpenStudy (***[isuru]***):

first of all there r only 2 choices which is given in point slope form... wt r they ?

OpenStudy (anonymous):

um A an B

OpenStudy (***[isuru]***):

for the point slope form...they give you a point (x1, y1) and a slope m, and have you plug it into this formula: y – y1 = m(x – x1) can u guess it now.. ?

OpenStudy (anonymous):

A an C i meant?

OpenStudy (***[isuru]***):

yep.. now according to the definition above.. can u guess the answer ?

OpenStudy (anonymous):

A?

OpenStudy (***[isuru]***):

(x1 , y1 ) = ( 6 , -3) which means x1 = 6 and y1 = -3 now the format is y – y1 = m(x – x1) wt is the answer then ?

OpenStudy (anonymous):

c

OpenStudy (***[isuru]***):

yep.. u r correct this time !!

OpenStudy (anonymous):

ok thx

OpenStudy (***[isuru]***):

u r welcome!